In triangle ABC, a = 6.8 , b = 15 , and sin C = 1/3. Find the area of triangle ABC.
Answer by Alan3354(69443)

In triangle ABC, a = 6.8 , b = 15 , and sin C = 1/3. Find the area of triangle ABC.
----------
Area = b*h/2
= 6.8*15*(1/3)/2
= 17 sq units